# SparseWorld-TC: Trajectory-Conditioned Sparse Occupancy World Model

### v2

@@ 1 Introduction
World models offer a principled framework for understanding environmental dynamics and are considered essential for AI systems operating in the physical world. Although the literature has yet to converge on a unified and explicit definition of world models, the ability to predict the evolution of the physical environment is consistently regarded as a core capability [suervey_1, suervey_2]. Among recent advances, occupancy-based world models have emerged as a prominent category [occsora, occworld, occ3d, occ-llm, i2world, dome, come], garnering significant interest due to their direct applicability in autonomous driving and robotics, as well as their role in generating temporally consistent sensor observations [dynamiccity, infinicube].
@@
Despite this progress, many recent approaches encode observations into discrete tokens using variational autoencoders (VAEs) [vae, vqvae] and then predict future occupancy from these limited-vocabulary representations [occworld, come, dome, i2world]. Such tokenization can constrain representational capacity and discard fine-grained information. Alternatively, some methods recast occupancy world modeling as predicting dynamic and static flow fields [uniocc, dfit-occworld], which does not inherently preserve generative capabilities.
Motivation. Pure attention-based feed-forward architectures have demonstrated remarkable flexibility and representational capacity in both language and 3D vision domains, as exemplified by models such as GPT [gpt4] and VGGT [vggt]. However, these capabilities remain underexplored for future occupancy prediction. Unlike methods that depend on fixed-resolution bird’s eye view (BEV) representations, which introduce explicit geometric constraints into feature interactions, we propose a pure attention-based transformer architecture that leverages fully sparse occupancy representations to directly capture spatiotemporal relationships from raw image features. Our approach eliminates both discretized tokens and explicit BEV geometric priors, enabling more effective integration of spatiotemporal dependencies, as illustrated in Figure 1.
Building upon recent advances in sparse perception models [sparsebev, sparsedrive], we represent scene occupancy using a set of anchors, each composed of a group of 3D query points and an associated feature vector. For each future frame, the anchor features interact through self-attention and cross-attention mechanisms with image features from past frames, accessed via deformable attention. This process iteratively refines the initially randomized points into accurate, context-aware occupancy representations. Trajectory information is incorporated as a conditioning signal to steer the temporal evolution of the predicted scene.
@@
  • Our method achieves state-of-the-art performance on the nuScenes benchmark [nuscenes, occ3d] for 1–3 second occupancy forecasting and demonstrates compelling results on long-term forecasting benchmarks [come, dome].
@@ 2 Related Work
For tokenization strategies, existing works [occworld, occsora, fsfnet, occllama, dome, renderworld, occ-llm, occvar, i2world, come] mainly adopt VAE [vae] or VQ-VAE [vqvae] based approaches. OccWorld [occworld] utilizes an occupancy VQ-VAE to discretize continuous 3D scene data into discrete tokens. OccLLaMA [occllama] extends this direction with a VAE-like scene tokenizer optimized for the sparsity and class imbalance of semantic occupancy, enabling unified visual-language-action modeling. RenderWorld [renderworld] introduces AM-VAE, a dual-codebook variant of VQ-VAE that separately encodes air and non-air voxels to enhance the granularity of the representation. I2-World [i2world] further advances VAE-based tokenization by designing a multi-scale architecture, which refines token granularity across different spatial scales to capture both local details and global scene structures. In contrast, DOME [dome] and COME [come] employ a continuous Occ-VAE instead of discrete tokenization, preserving fine-grained details through continuous latent spaces.
@@
Intermediate representations must balance precision and efficiency, including BEV feature [occworld, dfit-occworld, occvar, preworld, i2world, dome, come, occprophet, efficientocf], Gaussian representation [gaussianad, renderworld], and others like triplane [dtt] or hexplane [dynamiccity]. Most VAE-based methods [occworld, dome, come, occvar, i2world, occllama] rely on dense BEV feature maps for spatial-temporal modeling and conditional scene generation. PreWorld [preworld] validates BEV’s effectiveness through pre-training, while DFIT-OccWorld [dfit-occworld] uses it for flow prediction. GaussianAD [gaussianad] and RenderWorld [renderworld] adopt 3D Gaussian splatting [3dgaussian] to model scenes as anisotropic Gaussian ellipsoids, achieving efficient rendering and high segmentation accuracy. DTT [dtt] and DynamicCity [dynamiccity] utilize triplane or hexplane as intermediate representations, which retain 3D structural information while achieving more compact latent spaces.
@@
Regarding generation paradigms, autoregressive modeling [occworld, preworld, renderworld, occvar, i2world, drive-occworld, occ-llm, occllama] remains prevalent for sequential scene prediction, though non-autoregressive [dfit-occworld] and diffusion-based [dtt, dome, come] alternatives have gained traction. OccWorld [occworld] and RenderWorld [renderworld] use hierarchical transformers for autoregressive future scene token prediction. OccLLaMA [occllama] and Occ-LLM [occ-llm] leverage large language models to perform cross-modal next-token/scene autoregressive modeling. DFIT-OccWorld [dfit-occworld] stands out by adopting a non-autoregressive paradigm based on its BEV intermediate representations, which accelerates generation speed compared to autoregressive counterparts while maintaining scene integrity. DOME [dome] leverages a spatial-temporal diffusion transformer, and COME [come] extends this diffusion framework with ControlNet [controlnet] to enable conditional diffusion-based generation.
@@
Proposed Method. Rather than relying on handcrafted tokenizers or intermediate representations, our approach prefers to model the occupancy-based world in an end-to-end manner. Scenes are represented as sets of learnable feature embeddings, with their interactions mediated by attention mechanisms. This formulation departs from diffusion-based and autoregressive paradigms, favoring a VGGT-like feed-forward architecture [vggt] that predicts future occupancy in a single forward pass. Under this design, conditioning scene evolution on a given trajectory is straightforward: we embed the trajectory as feature vectors and let them interact with the scene features.
@@ 3.1 Sparse World Model Representation
Most existing approaches model 3D scene occupancy using dense voxel grids [fbocc], which are typically constructed on BEV representations. Although effective, such BEV-based paradigms impose explicit geometric constraints that can limit the flexible interaction of spatiotemporal features across different scales. Inspired by recent advances in sparse perception models that bypass BEV as an intermediate representation [petrv2, sparse4d, sparsebev, opus], we propose an alternative occupancy formulation based on a collection of anchors. In our framework, each anchor consists of a set of randomly initialized 3D points and an associated feature vector that jointly predicts per-point geometric offsets and semantic labels, as illustrated in Figure 2. Based on this sparse occupancy representation, we define the world model as a sequence of such single-frame occupancy states conditioned on a given trajectory.
@@ 3.2.1 Trajectory Spatiotemporal Embedding
Position Embedding. In trajectory representation, position embedding captures the spatial characteristics of trajectories by integrating the spatial information of each waypoint into the feature system. Specifically, the 3D coordinates of waypoints first undergo a relative pose transformation for subsequent fusion, then are mapped to the target feature dimension via a multilayer perceptron (MLP). The homogeneous matrix of the relative pose transformation is also mapped to the feature dimension using an MLP, ultimately equipping the features with spatial attributes and ego trajectory information to support subsequent trajectory-conditioned occupancy forecasting.
Time Embedding. Time embedding captures the relative temporal positional relationship of trajectories using a list of future timestamps (e.g. [0, 2, 4, 6]), where ”0” denotes the current frame timestamp and other values represent the number of frames relative to the current frame. Classic sine-cosine encoding is employed to embed this relative temporal relationship into the corresponding features.
Spatiotemporal Embedding. Drawing inspiration from motion-aware layer normalization (MLN) [streampetr], spatiotemporal embedding achieves the fusion of spatial and temporal information by building on the aforementioned position embedding and time embedding. Specifically, two linear layers are designed to implicitly learn the affine transformations between adjacent frames. This learning process relies on the spatial attribute cues provided by position embedding and the relative temporal relationship information captured by time embedding, enabling the affine transformations to adapt dynamically to the spatiotemporal characteristics of the trajectory. Subsequently, the implicitly learned affine transformations are applied to the query features: by adjusting the query features through these transformations, the spatial information and temporal information are organically integrated into the query features. The final trajectory $\tau$ is thus represented as a set of spatiotemporal feature vectors $\tau=\{\mathbf{f}^{\prime\prime}_{t}\}_{t=1}^{T}$, where $\mathbf{f}^{\prime\prime}_{t}$ denotes the spatiotemporal feature embedding corresponding to the $t$ frame.
@@ 3.2.2 Sensor Embedding with Deformable Attention
Deformable attention is widely used in 3D sparse perception [detr3d, bevformer, sparse4d, sparsebev, sparsedrive]. It projects 3D query points into multi-view image feature maps (and 3D LiDAR voxel grids when LiDAR is available) using known calibration and ego-motion, and samples features at the projected locations.
In our architecture, we take into account $m$ sensors per frame, the center $\mathbf{c}_{i}$ of each anchor point set $P_{i}$ serves as the query point in the deformable attention operation, where we calculate the mean and standard deviation of anchor point set $P_{i}$ along the $x,y,z$ directions as the basis for sampling offsets. Each center is projected into the multi-scale image feature maps of the backbone (e.g. ResNet [resnet] or ViT [vit]) using camera intrinsics, extrinsics and the ego pose. If a query projects into multiple views due to field-of-view overlap, we aggregate by averaging the sampled features across all $m$ views. Each anchor center collects features from the past $T^{\prime}$ frames. To encode temporal context, we add a sinusoidal time embedding processed by a fully connected layer to provide motion cues. Finally, for each future frame with occupancy representation $\mathcal{M}$, we obtain the sensor embedding representation $\mathcal{S}=\{\mathbf{f}^{\prime}_{i}\}_{i=1}^{N\times T^{\prime}}$.
@@ 3.2.3 Fully Attention Fusion Architecture
Table 1: 4D occupancy forecasting performance on Occ3D-nuScenes [occ3d]. Input denotes the modality of input apart from the ego trajectory. Avg. denotes average performance of that in 1s, 2s, and 3s.
@@
OccWorld-O [occworld] Occ GT 25.78 15.14 10.51 17.14 34.63 25.07 20.18 26.63 18.00
OccLLaMA-O [occllama] Occ GT 25.05 19.49 15.26 19.93 34.56 28.53 24.41 29.17 -
RenderWorld [renderworld] Occ GT 28.69 18.89 14.83 20.89 37.74 28.41 24.08 30.08 -
Occ-LLM [occ-llm] Occ GT 24.02 21.65 17.29 20.99 36.65 32.14 28.77 35.52 -
DFIT-OccWorld [dfit-occworld] Occ GT 31.68 21.29 15.19 22.71 40.28 31.24 25.29 32.27 -
DOME [dome] Occ GT 35.11 25.89 20.29 27.10 43.99 35.36 29.74 36.36 6.54
UniScene [uniscene] Occ GT 35.37 29.59 25.08 31.76 38.34 32.70 29.09 34.84 -
I2-World [i2world] Occ GT 47.62 38.58 32.98 39.73 54.29 49.43 45.69 49.80 37.04
RenderWorld [renderworld] Camera 2.83 2.55 2.37 2.58 14.61 13.61 12.98 13.73 -
OccWorld-D [occworld] Camera 11.55 8.10 6.22 8.62 18.90 16.26 14.43 16.53 -
OccLLaMA [occllama] Camera 10.34 8.66 6.98 8.66 25.81 23.19 19.97 22.99 -
PreWorld [preworld] Camera 12.27 9.24 7.15 9.55 23.62 21.62 19.63 21.62 -
Occ-LLM [occ-llm] Camera 11.28 10.21 9.13 10.21 27.11 24.07 20.19 23.79 -
DFIT-OccWorld [dfit-occworld] Camera 13.38 10.16 7.96 10.50 19.18 16.85 25.02 17.02 -
OccVAR [occvar] Camera 17.17 10.38 7.82 11.79 27.60 25.14 20.33 24.35 -
SparseWorld [dang2025sparseworld] Camera 14.93 13.15 11.51 13.20 22.96 22.10 21.05 22.03 8.0
DOME-STC [dome] Camera 17.79 14.23 11.58 14.53 26.39 23.20 20.42 23.33 2.75
I2-World-STC [i2world] Camera 21.67 18.78 16.47 18.97 30.55 28.76 26.99 28.77 4.21
DOME-F [dome] Camera 24.12 17.41 13.24 18.25 35.18 27.90 23.44 28.84 -
DTT [dtt] Camera 24.87 18.30 15.63 19.60 38.98 37.45 31.89 36.11 -
COME-F [come] Camera 26.56 21.73 18.49 22.26 48.08 43.84 40.28 44.07 -
@@ 3.3 Training Strategies
Loss Term. We extract the center of each ground-truth occupancy voxel as target points and optimize the Chamfer Distance loss $\mathcal{L}_{CD}$ (Equation 5) to align the predicted point distribution with the target points. This loss function is widely adopted in point cloud processing [khurana2023point, yang2024visual] and occupancy modeling [opus], as it effectively measures the similarity between predicted point clouds $\mathbb{P}$ and target point clouds $\mathbb{P}_{g}$. This loss thereby supervises the predicted offsets, guiding the initially randomized points toward the precise underlying geometry.

### v3

@@ 1 Introduction
World models offer a principled framework for understanding environmental dynamics and are considered essential for AI systems operating in the physical world. Although the literature has yet to converge on a unified and explicit definition of world models, the ability to predict the evolution of the physical environment is consistently regarded as a core capability [17, 6, 19]. Among recent advances, occupancy-based world models have emerged as a prominent category [42, 55, 38, 49, 25, 8, 34], garnering significant interest due to their direct applicability in autonomous driving and robotics, as well as their role in generating temporally consistent sensor observations [2, 32].
@@
Despite this progress, many recent approaches encode observations into discrete tokens using variational autoencoders (VAEs) [16, 39] and then predict future occupancy from these limited-vocabulary representations [55, 34, 8, 25]. Such tokenization can constrain representational capacity and discard fine-grained information. Alternatively, some methods recast occupancy world modeling as predicting dynamic and static flow fields [45, 53], which does not inherently preserve generative capabilities.
Motivation. Pure attention-based feed-forward architectures have demonstrated remarkable flexibility and representational capacity in both language and 3D vision domains, as exemplified by models such as GPT [1] and VGGT [41]. However, these capabilities remain underexplored for future occupancy prediction. Unlike methods that depend on fixed-resolution bird’s eye view (BEV) representations, which introduce explicit geometric constraints into feature interactions, we propose a pure attention-based transformer architecture that leverages fully sparse occupancy representations to directly capture spatiotemporal relationships from raw image features. Our approach eliminates both discretized tokens and explicit BEV geometric priors, enabling more effective integration of spatiotemporal dependencies, as illustrated in Figure 1.
Building upon recent advances in sparse perception models [28, 36], we represent scene occupancy using a set of anchors, each composed of a group of 3D query points and an associated feature vector. For each future frame, the anchor features interact through self-attention and cross-attention mechanisms with image features from past frames, accessed via deformable attention. This process iteratively refines the initially randomized points into accurate, context-aware occupancy representations. Trajectory information is incorporated as a conditioning signal to steer the temporal evolution of the predicted scene.
@@
  • Our method achieves state-of-the-art performance on the nuScenes benchmark [3, 38] for 1–3 second occupancy forecasting and demonstrates compelling results on long-term forecasting benchmarks [34, 8].
@@ 2 Related Work
For tokenization strategies, existing works [55, 42, 9, 46, 8, 50, 49, 12, 25, 34] mainly adopt VAE [16] or VQ-VAE [39] based approaches. OccWorld [55] utilizes an occupancy VQ-VAE to discretize continuous 3D scene data into discrete tokens. OccLLaMA [46] extends this direction with a VAE-like scene tokenizer optimized for the sparsity and class imbalance of semantic occupancy, enabling unified visual-language-action modeling. RenderWorld [50] introduces AM-VAE, a dual-codebook variant of VQ-VAE that separately encodes air and non-air voxels to enhance the granularity of the representation. I2-World [25] further advances VAE-based tokenization by designing a multi-scale architecture, which refines token granularity across different spatial scales to capture both local details and global scene structures. In contrast, DOME [8] and COME [34] employ a continuous Occ-VAE instead of discrete tokenization, preserving fine-grained details through continuous latent spaces.
@@
Intermediate representations must balance precision and efficiency, including BEV feature [55, 53, 12, 20, 25, 8, 34, 4, 48], Gaussian representation [56, 50], and others like triplane [47] or hexplane [2]. Most VAE-based methods [55, 8, 34, 12, 25, 46] rely on dense BEV feature maps for spatial-temporal modeling and conditional scene generation. PreWorld [20] validates BEV’s effectiveness through pre-training, while DFIT-OccWorld [53] uses it for flow prediction. GaussianAD [56] and RenderWorld [50] adopt 3D Gaussian splatting [14] to model scenes as anisotropic Gaussian ellipsoids, achieving efficient rendering and high segmentation accuracy. DTT [47] and DynamicCity [2] utilize triplane or hexplane as intermediate representations, which retain 3D structural information while achieving more compact latent spaces.
@@
Regarding generation paradigms, autoregressive modeling [55, 20, 50, 12, 25, 51, 49, 46] remains prevalent for sequential scene prediction, though non-autoregressive [53] and diffusion-based [47, 8, 34] alternatives have gained traction. OccWorld [55] and RenderWorld [50] use hierarchical transformers for autoregressive future scene token prediction. OccLLaMA [46] and Occ-LLM [49] leverage large language models to perform cross-modal next-token/scene autoregressive modeling. DFIT-OccWorld [53] stands out by adopting a non-autoregressive paradigm based on its BEV intermediate representations, which accelerates generation speed compared to autoregressive counterparts while maintaining scene integrity. DOME [8] leverages a spatial-temporal diffusion transformer, and COME [34] extends this diffusion framework with ControlNet [54] to enable conditional diffusion-based generation.
@@
Proposed Method. Rather than relying on handcrafted tokenizers or intermediate representations, our approach prefers to model the occupancy-based world in an end-to-end manner. Scenes are represented as sets of learnable feature embeddings, with their interactions mediated by attention mechanisms. This formulation departs from diffusion-based and autoregressive paradigms, favoring a VGGT-like feed-forward architecture [41] that predicts future occupancy in a single forward pass. Under this design, conditioning scene evolution on a given trajectory is straightforward: we embed the trajectory as feature vectors and let them interact with the scene features.
@@ 3.1 Sparse World Model Representation
Most existing approaches model 3D scene occupancy using dense voxel grids [21], which are typically constructed on BEV representations. Although effective, such BEV-based paradigms impose explicit geometric constraints that can limit the flexible interaction of spatiotemporal features across different scales. Inspired by recent advances in sparse perception models that bypass BEV as an intermediate representation [29, 27, 28, 40], we propose an alternative occupancy formulation based on a collection of anchors. In our framework, each anchor consists of a set of randomly initialized 3D points and an associated feature vector that jointly predicts per-point geometric offsets and semantic labels, as illustrated in Figure 2. Based on this sparse occupancy representation, we define the world model as a sequence of such single-frame occupancy states conditioned on a given trajectory.
@@ 3.2.1 Trajectory Spatiotemporal Embedding
Position Embedding. Each waypoint is defined by a 7-dimensional global coordinate $(x,y,z,q_{0},q_{1},q_{2},q_{3})$ and transformed to a 16-dimensional homogeneous matrix (relative coordinate in the current ego). The positional feature vector $\mathbf{f}_{i}^{P_{n}}$ is obtained by an MLP that projects the 16-dimensional matrix to the same dimension as the corresponding feature vector $\mathbf{f}_{i}$ for occupancy anchor $i$.
Time Embedding. Under the assumption of a fixed raw sensor frequency, the timestamps can be represented as $0,1,\dots,n]$, where 0 denotes the current timestamp. For timestamp $n$, this integer is encoded into a positional embedding $\mathbf{f}_{i}^{T_{n}}$ using standard sinusoidal positional encoding.
Spatiotemporal Embedding. Drawing inspiration from motion-aware layer normalization (MLN) [43], spatiotemporal embedding achieves the fusion of spatial and temporal information by building on the aforementioned position embedding and time embedding. Specifically, two linear layers are designed to implicitly learn the affine transformations between adjacent frames. The spatiotemporal embedding method is formulated as: $\mathbf{f}^{\prime\prime}_{t}=\gamma(\mathbf{f}_{i}^{P_{n}})\cdot(\mathbf{f}_{i}+\mathbf{f}_{i}^{T_{n}})+\beta(\mathbf{f}_{i}^{P_{n}})$, where both $\gamma$ and $\beta$ are distinct linear layers applied to $\mathbf{f}_{i}^{P_{n}}$. The final trajectory $\tau$ is thus represented as a set of spatiotemporal feature vectors $\tau=\{\mathbf{f}^{\prime\prime}_{t}\}_{t=1}^{T}$.
@@ 3.2.2 Sensor Embedding with Deformable Attention
Deformable attention is widely used in 3D sparse perception [44, 22, 27, 28, 36]. It projects 3D query points into multi-view image feature maps (and 3D LiDAR voxel grids when LiDAR is available) using known calibration and ego-motion, and samples features at the projected locations.
In our architecture, we take into account $m$ sensors per frame, the center $\mathbf{c}_{i}$ of each anchor point set $P_{i}$ serves as the query point in the deformable attention operation, where we calculate the mean and standard deviation of anchor point set $P_{i}$ along the $x,y,z$ directions as the basis for sampling offsets. Each center is projected into the multi-scale image feature maps of the backbone (e.g. ResNet [10] or ViT [7]) using camera intrinsics, extrinsics and the ego pose. If a query projects into multiple views due to field-of-view overlap, we aggregate by averaging the sampled features across all $m$ views. Each anchor center collects features from the past $T^{\prime}$ frames. To encode temporal context, we add a sinusoidal time embedding processed by a fully connected layer to provide motion cues. Finally, for each future frame with occupancy representation $\mathcal{M}$, we obtain the sensor embedding representation $\mathcal{S}=\{\mathbf{f}^{\prime}_{i}\}_{i=1}^{N\times T^{\prime}}$.
@@ 3.2.3 Fully Attention Fusion Architecture
Table 1: 4D occupancy forecasting performance on Occ3D-nuScenes [38]. Input denotes the modality of input apart from the ego trajectory. Avg. denotes average performance of that in 1s, 2s, and 3s.
@@
OccWorld-O [55] Occ GT 25.78 15.14 10.51 17.14 34.63 25.07 20.18 26.63 18.00
OccLLaMA-O [46] Occ GT 25.05 19.49 15.26 19.93 34.56 28.53 24.41 29.17 -
RenderWorld [50] Occ GT 28.69 18.89 14.83 20.89 37.74 28.41 24.08 30.08 -
Occ-LLM [49] Occ GT 24.02 21.65 17.29 20.99 36.65 32.14 28.77 35.52 -
DFIT-OccWorld [53] Occ GT 31.68 21.29 15.19 22.71 40.28 31.24 25.29 32.27 -
DOME [8] Occ GT 35.11 25.89 20.29 27.10 43.99 35.36 29.74 36.36 6.54
UniScene [18] Occ GT 35.37 29.59 25.08 31.76 38.34 32.70 29.09 34.84 -
I2-World [25] Occ GT 47.62 38.58 32.98 39.73 54.29 49.43 45.69 49.80 37.04
RenderWorld [50] Camera 2.83 2.55 2.37 2.58 14.61 13.61 12.98 13.73 -
OccWorld-D [55] Camera 11.55 8.10 6.22 8.62 18.90 16.26 14.43 16.53 -
OccLLaMA [46] Camera 10.34 8.66 6.98 8.66 25.81 23.19 19.97 22.99 -
PreWorld [20] Camera 12.27 9.24 7.15 9.55 23.62 21.62 19.63 21.62 -
Occ-LLM [49] Camera 11.28 10.21 9.13 10.21 27.11 24.07 20.19 23.79 -
DFIT-OccWorld [53] Camera 13.38 10.16 7.96 10.50 19.18 16.85 25.02 17.02 -
OccTENS-F [13] Camera 17.17 10.38 7.82 11.79 27.60 25.14 20.33 24.35 -
SparseWorld [5] Camera 14.93 13.15 11.51 13.20 22.96 22.10 21.05 22.03 8.00
DOME-STC [8] Camera 17.79 14.23 11.58 14.53 26.39 23.20 20.42 23.33 2.75
I2-World-STC [25] Camera 21.67 18.78 16.47 18.97 30.55 28.76 26.99 28.77 4.21
DOME-F [8] Camera 24.12 17.41 13.24 18.25 35.18 27.90 23.44 28.84 -
DTT [47] Camera 24.87 18.30 15.63 19.60 38.98 37.45 31.89 36.11 -
COME [34] Camera 26.56 21.73 18.49 22.26 48.08 43.84 40.28 44.07 -
@@ 3.3 Training Strategies
Loss Term. We extract the center of each ground-truth occupancy voxel as target points and optimize the Chamfer Distance loss $\mathcal{L}_{CD}$ (Equation 5) to align the predicted point distribution with the target points. This loss function is widely adopted in point cloud processing [15, 52] and occupancy modeling [40], as it effectively measures the similarity between predicted point clouds $\mathbb{P}$ and target point clouds $\mathbb{P}_{g}$. This loss thereby supervises the predicted offsets, guiding the initially randomized points toward the precise underlying geometry.